OpenStudy (anonymous):

Which equation will result if the substitution method is used to solve the following system of linear equations? x – y = 7 x = y + 3 A. (y + 3) – y = 7 B. x – y = y + 3 C. x = (7 – x) + 3 D. x – (y + 3) = 7

OpenStudy (anonymous):

If you help me quickly, I will fan and medal. This is my last question! :)

OpenStudy (anonymous):

It is nice that one of those equations has a variable alone on one side. The second equation can be used to substitute the X in the first equation with (y+3). Just make that substitution and you will have your answer.

OpenStudy (anonymous):

I am terrible at this. I don't understand...:/

OpenStudy (anonymous):

That's okay, we just have to go about it a different way :) Okay, so you have: x - y = 7 x = y + 3 Now, see that in Equation 2, it states that X = y + 3. Well, now that you know that X = (y + 3), you can plug (y + 3) into anywhere that X appears in the other equation! :D Soooo, look at Equation 1 again. See any X's? Now just replace (or "substitute") that X with your (Y + 3) :) Let me know if you're still confused so I can help!

OpenStudy (johnweldon1993):

Well notice x – y = 7 x = y + 3 That second equation there...tells us what 'x' equals.....x equals (y + 3) since we know that...we can use it in the first equation....now our first equation was x - y = 7 But since we know that x = y + 3 we can now replace every 'x' we see ..with what 'x' equals...so for every 'x' you see...replace it with the y + 3 (y + 3) - y = 7